OpenStudy (anonymous):

Help? Linear equations using elimination? The perimeter of a rectangle is 48 cm. The length is 6 cm less than twice the width. Which system would model this situation, and what are the length and the width? A. 2l + 2w = 48 l + 2w = 6 Solution: (9, 6) B. 2l + 2w = 48 2l + w = –6 Solution: (–32, 56) C. l + w = 48 l – 2w = –6 Solution: (18, 6) D. 2l + 2w = 48 l – 2w = –6 Solution: (14, 10)

OpenStudy (anonymous):

First you have to understand the problem. What formula do you need in order to model the problem?

OpenStudy (anonymous):

The perimeter of a rectangle which is: \(P=2l+2w\)

OpenStudy (anonymous):

Then write down all the given: \(\Large {P=48 cm,\\ l=2w-6} \)

OpenStudy (anonymous):

After these steps, substitute all the given into the formula

OpenStudy (anonymous):

Tell me what you got and i'll check it later :)

OpenStudy (anonymous):

I still don't understand... :(

OpenStudy (anonymous):

which part you don't understand? O.o

OpenStudy (anonymous):

I understand it, I just don't know where to substitute and what to substitute in.... Because I have to use elimination first..

OpenStudy (anonymous):

you can do it later after you find all the equation you need. After that, you can use elimination to find the solutions. so let's do substitution to find the first equation :) since P=2l+2w and P=48, you will have \(2l+2w=48\)

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

Okay? There are two variables and you can't necessarily solve an equation once you've used substitution. I'm just so lost.

OpenStudy (anonymous):

You just have to substitute the given into the formula for the perimeter of the rectangle to get an \(equation\). Think about this, how can you apply elimination if you don't have equations? so you already have : 2l +2w = 48 as your first equation From the given, you can find the second equation which is \(l=2w−6\) you can rearrange this equation so that they will have the same format as the first equation: \(l-2w=6\)

OpenStudy (anonymous):

i made a mistake, it should be: \(\LARGE l−2w=-6\), i forgot the negative.

OpenStudy (anonymous):

did you understand already? I just used substitutions to get equations then you can apply elimination later once you have the equations :)

OpenStudy (anonymous):

I'll try and solve it. Thank you.

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

Wait, so how am I supposed to use elimination when I can't solve this equation using elimination. \[21+2w=48\] \[l-2w=-6\]

OpenStudy (anonymous):

For example, if I had 3y-4x=9 and -3y+8x=-2 I could easily use elimination because both equations have a y and x variable.

OpenStudy (anonymous):

you can. just add the two equations and the variable w will be eliminated then solve for l.

OpenStudy (anonymous):

it's the same thing. instead of x and y, you are using l and w in this problem.

OpenStudy (anonymous):

Okayyy I seeeee Thank youuu :D

OpenStudy (anonymous):

Is the solution (14,10)?

OpenStudy (anonymous):

great job! :) it's right

OpenStudy (anonymous):

Thank you for helping me.

OpenStudy (anonymous):

you're welcome as long as you understand it there will be no problem ;) goodluck!